OpenStudy (anonymous):

convert 120 degrees to radians in terms of pi

OpenStudy (anonymous):

do you have a unit circle?

hartnn (hartnn):

to convert degrees to radians just multiply it by \(\large \dfrac{\pi}{180}\)

hartnn (hartnn):

\(\Large x^r =x^0 \dfrac{\pi}{180} \)

hartnn (hartnn):

your \(x^o =120\) here

OpenStudy (anonymous):

@hartnn is right, you just simplify from there.

OpenStudy (perl):